OBSERVATION AND ANALYSIS OF THE ν_{3} and 2ν_{3} BANDS OF ^{32}S^{18}O_{3} AND ^{34}S^{18}O_{3}

STEVEN W. SHARPE, ROBERT L. SAMS, \underline{THOMAS A. BLAKE}, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC06-76RLO 1830); ARTHUR MAKI, 15012 24th Ave. S. E. Mill Creek, WA 98012; NICO-LAE VULPANOVICI, TONY MASIELLO, JEFFREY BARBER, ENGELENE t.H. CHRYSOS-TOM, JOSEPH W. NIBLER, Department of Chemistry, Oregon State University, 153 Gilbert Hall, Corvallis, OR 97331-4003; ALFONS WEBER, National Institute of Standards and Technology, Gaithersburg, MD 20899 and National Science Foundation, 4201 Wilson Boulevard, Arlington, VA 22230.

We are engaged in a comprehensive investigation of the fundamental and low lying combination and overtone bands of the parent and some isotopic variants of sulfur trioxide. Results and analysis for the Raman^{a} and infrared^{b} bands of the parent isotopomer have been published recently. Here, results and analysis of the infrared absorption spectra of the ν_{3} and 2ν_{3} bands of ^{32}S^{18}O_{3} and ^{34}S^{18}O_{3} recorded separately at a resolution of 0.002 cm^{−1}will be presented. For ^{32}S^{18}O_{3}, ν_{3} is centered at 1348.85320(3) cm^{−1}with B′ = 0.30879466(7) cm^{−1}, C′ = 0.15436012(6) cm^{−1}and C′ζ = 0.0795934(4) cm^{−1}. The l = ±2 component of the first overtone of ν_{3} is centered at 2692.8899(3) cm^{−1}with B′ = 0.307815(1) cm^{−1}, C′ = 0.1538331(7) cm^{−1}and C′ζ = 0.078716(2) cm^{−1}. For ^{34}S^{18}O_{3}, ν_{3} is centered at 1330.07810(1) cm^{−1}with B′ = 0.3088233(3) cm^{−1}and C′ = 0.15437917(3) cm^{−1}, and C′ζ = 0.0770680(3) cm^{−1}. The l = ±2 component of the first overtone of ν_{3} is centered at 2655.45735(4) cm^{−1}with B′ = 0.3078634(8) cm^{−1}, C′ = 0.15386494(7) cm^{−1}and C′ζ = 0.0762487(3) cm^{−1}. The Hamiltonian for analyzing these bands including splitting and perturbing terms and other fit constants will be presented.

aE. t.H. Chrysostom, N. Vulpanovici, T. Masiello, J. Barber, J. W. Nibler, A. Weber, A. Maki, T. A. Blake, J. Mol. Spectrosc.

210, 233 (2001).

bA. Maki, T. A. Blake, R. L. Sams, N. Vulpanovici, J. Barber, E. t.H. Chrysostom, T. Masiello, J. W. Nibler, A. Weber, J. Mol. Spectrosc. 210, 240 (2001).